# Causal Inference in Longitudinal Data under Unknown Interference

### v4

@@ 6 Application
In this section, we illustrate the application of the method using one empirical example in a spatial setting and the other in a social network setting. Simulation results on the method’s performance are reported in the Supplementary Material, which show that the WLS estimator yields consistent estimates and confidence intervals with desirable coverage rates.
@@ 6.1 Public Project and Political Support
We first apply our method to the study in Stokes (2016) on the political consequences of building wind turbines in Ontario, Canada, where the Liberal Party government passed laws since 2003 enabling individuals and corporations to build wind turbines, despite objections from local communities. The study collected election data from 6,186 precincts in Ontario over three elections: 2003, 2007, and 2011, and combined the data with the location of each proposed turbine. Precinct $i$ is considered as treated in year $t$ if a turbine project within its boundary has been proposed before the election. The data exhibit the structure of staggered adoption, where once a unit is treated, it remains treated for the remaining periods. The three treatment histories from 2007 to 2011 ($t=1$ and $t=2$) are $(0,0)$, $(0,1)$, $(1,1)$. There were 6,002 never treated precincts, 53 treated since 2007, and 131 treated since 2011.111111The original analysis combines the difference-in-differences estimator with an exposure mapping, hence its validity hinges on the correct model specification. The outcomes of interest are turnout rate and the Liberal Party’s vote share in 2011.
@@
The results are presented in the top panel of Figure 3. On the left side are the estimated causal effects for the turnout rate, and on the right side are those for the Liberal Party’s vote share. The black dots in the figures represent estimates for $\mu((1,1);d)-\mu((0,0);d)$, and the gray dots represent estimates for $\mu((0,1);d)-\mu((0,0);d)$. We use superscripts on $\mu(\cdot)$ to indicate the time at which the outcome is measured. The segments denote the $95\%$ confidence intervals of these estimates. The estimates provide evidence for the presence of spillover effects generated by both histories. Being treated in both $2007$ and $2011$ would increase the turnout rate by $7.7\%$ and decrease the Liberal Party’s vote share by $6.4\%$ in a precinct. The effects are also observed among its neighbors that are $2$-$4$ km away, where the turnout rate would increase by $5.0\%$, and the Liberal Party’s vote share would decrease by 6.0%. In precincts treated only in $2011$, treatment assignment history caused a rise of $3.1\%$ in the turnout rate and a decline of $7.0\%$ in the Liberal Party’s vote share. The spillover effects on its neighbors that are $2$-$4$ km away are of similar magnitude. The effects on the outcome variables become close to zero (and statistically insignificant) among precincts that are more than $6$ km away from the treatment, which justifies the choice of the cutoff values.
@@ 6.2 Impacts of Smoking Cessation on Health Outcomes
Our second application examines how quitting smoking affects both an individual’s health and that of their social network neighbors. Prior research has established strong links between smoking and adverse health outcomes (Jain et al., 2016). It is natural to conjecture that reducing exposure to second-hand smoke could yield benefits for those nearby. We use data from the Framingham Heart Study (FHS), focusing on the offspring cohort and variables from the first three exams. The average time gap is 8 years between exams 1 and 2, and 5 years between exams 2 and 3. We treat exam 1 as the initial period ($t=0$) and define treatment history using exams 2 and 3 ($t\in\{1,2\}$). $A_{i}^{(t)}=1$ if the individual reports not smoking at time $t$. Outcomes include body weight, blood pressure, cholesterol level, and incidence of cardiovascular disease (CVD). The first three are measured at exam 3 ($t=2$); CVD incidence is defined by the occurrence of symptoms (e.g., stroke) within 10 years after exam 3.121212We examine whether there is any CVD incidence between exam 4 and exam 6. The attrition rate during this period is 16.4%, therefore the influence of righ-censoring is limited. FHS records social ties across multiple dimensions (e.g., spouse, family, friends, neighbors, etc.). We define two individuals as connected if any type of social tie exists between them and include only ties formed prior to exam 2 to avoid endogeneity in network formation.
@@
Additional results from both applications are provided in the Supplementary Material, including placebo tests supporting the identification assumptions, estimates based on a propensity score model that accounts for treatment diffusion, and estimated effects on blood pressure and cholesterol levels in the FHS.
@@ 7 Conclusion
This paper tackles the challenge of causal inference in longitudinal data with an unknown interference structure. We introduce a novel estimand, the AMR, to capture both direct and spillover effects generated by different treatment histories. This estimand is well-defined whenever a proximity metric is available, requiring only minimal knowledge of the underlying interference structure. It can be summarized by an MSM and identified under a slightly generalized form of sequential exchangeability. We show how to estimate any AMR and MSM parameters using a WLS estimator and establish its large-sample properties, including consistency and asymptotic normality, under mild restrictions on the dependence across units. We also develop HAC variance estimators for spillover effect estimates that are ensured to be asymptotically conservative. Our approach bridges methods in longitudinal data analysis and causal inference under interference, providing researchers with a powerful tool to study a broad range of causal effects in spatial and network settings, especially when interactions among units are poorly understood. We illustrate the method’s utility through simulations and two empirical applications, and conclude by discussing possible extensions.
Compared to alternative approaches based on outcome models or exposure mappings, our method avoids risks of model misspecification and the burden of computing exposure probabilities in settings with complex and unknown interference. However, it relies on researchers’ substantive understanding of the treatment assignment mechanism so that the propensity score is either known or can be consistently estimated. Future research should investigate more flexible strategies for propensity score estimation, such as sieve estimators (Hirano, Imbens and Ridder, 2003) or highly adaptive LASSO (Ertefaie, Hejazi and van der Laan, 2020). Another promising direction is to combine our approach with outcome modeling to improve robustness and efficiency.
@@
Appendix A Proofs
@@ A.1 Structural Equations under the Identification Assumptions
Here, $\varepsilon_{\mathbf{L}i}^{(t)}$, $\varepsilon_{Ai}^{(t)}$, and $\varepsilon_{Yi}$ represent other factors that may influence $\mathbf{L}_{i}^{(t)}$, $A_{i}^{(t)}$, and $Y_{i}$, respectively. They may include a unit’s own attributes, characteristics of its neighbors, and measurement errors. We do not assume that these error terms are identically distributed. $\mathcal{N}_{\mathcal{G}_{N}}(i)$ in the subscripts denotes a unit-specific and unknown neighborhood for unit $i$ in $\mathcal{G}_{N}$. The equations are compatible with the DAG in Figure 1 of the main text and closely resemble those in Ogburn et al. (2020). However, our framework allows for arbitrary heterogeneity in how units respond to treatment histories and does not impose a predefined exposure mapping that summarizes the effects of $\mathbf{A}_{\mathcal{N}}^{(1:T)}$. These equations formally define what unknown interference means and reflect a more realistic setting for spatial or network data, where both $Y_{i}$ and $\mathbf{L}_{i}^{(t)}$ may be shaped by unobserved and idiosyncratic factors, such as unit-specific neighborhood structures. This flexibility introduces the additional challenges for inference discussed below, similar to those encountered in design-based frameworks, although we allow for stochastic variation beyond the treatment assignment mechanism.
A.2 Proof for Proposition 1

### v5

@@ 6 Application
In this section, we illustrate the application of the method using one empirical example in a spatial setting and the other in a social network setting. Simulation results on the method’s performance are reported in Section C.2 of the Supplementary Material, which show that the WLS estimator generates consistent estimates and confidence intervals with desirable coverage rates.
@@ 6.1 Public Project and Political Support
We first apply our method to the study in Stokes (2016) on the political consequences of building wind turbines in Ontario, Canada, where the Liberal Party government passed laws since 2003 enabling individuals and corporations to build wind turbines, despite objections from local communities. The study collected election data from 6,186 precincts in Ontario over three elections: 2003, 2007, and 2011, and combined the data with the location of each proposed turbine. Precinct $i$ is considered as treated in year $t$ if a turbine project within its boundary has been proposed before the election. The data exhibit the structure of staggered adoption, where once a unit is treated, it remains treated for the remaining periods. The three treatment histories from 2007 to 2011 ($t=1$ and $t=2$) are $(0,0)$, $(0,1)$, $(1,1)$. There were 6,002 never treated precincts, 53 treated since 2007, and 131 treated since 2011.161616The original analysis combines the difference-in-differences estimator with an exposure mapping, hence its validity hinges on the correct model specification. The outcomes of interest are turnout rate and the Liberal Party’s vote share in 2011.
@@
The results are presented in the top panel of Figure 3. On the left side are the estimated causal effects for the turnout rate, and on the right side are those for the Liberal Party’s vote share. The black dots in the figures represent estimates for $\mu((1,1);d)-\mu((0,0);d)$, and the gray dots represent estimates for $\mu((0,1);d)-\mu((0,0);d)$. We use superscripts on $\mu(\cdot)$ to indicate the time at which the outcome is measured. The segments denote the $95\%$ confidence intervals of these estimates. The estimates provide evidence for the presence of spillover effects generated by both histories. Being treated in both $2007$ and $2011$ would increase the turnout rate by $7.7\%$ and decrease the Liberal Party’s vote share by $6.4\%$ in a precinct. The effects are also observed among its neighbors that are 22-44 km away, where the turnout rate would increase by $5.0\%$, and the Liberal Party’s vote share would decrease by 6.0%. In precincts treated only in $2011$, treatment assignment history caused a rise of $3.1\%$ in the turnout rate and a decline of $7.0\%$ in the Liberal Party’s vote share. The spillover effects on its neighbors that are 22-44 km away are of similar magnitude. The effects on the outcome variables become close to zero (and statistically insignificant) among precincts that are more than $6$ km away from the treatment, which justifies the choice of the cutoff values.
@@ 6.2 Impacts of Smoking Cessation on Health Outcomes
Our second application examines how quitting smoking affects both an individual’s health and that of their social network neighbors. Prior research has established strong links between smoking and adverse health outcomes (Jain et al., 2016). It is natural to conjecture that reducing exposure to second-hand smoke could yield benefits for those nearby. We use data from the Framingham Heart Study (FHS), focusing on the offspring cohort and variables from the first three exams. The average time gap is 8 years between exams 1 and 2, and 5 years between exams 2 and 3. We treat exam 1 as the initial period ($t=0$) and define treatment history using exams 2 and 3 ($t\in\{1,2\}$). $A_{i}^{(t)}=1$ if the individual reports not smoking at time $t$. Outcomes include body weight, blood pressure, cholesterol level, and incidence of cardiovascular disease (CVD). The first three are measured at exam 3 ($t=2$); CVD incidence is defined by the occurrence of symptoms (e.g., stroke) within 10 years after exam 3.171717We examine whether there is any CVD incidence between exam 4 and exam 6. The attrition rate during this period is 16.4%, therefore the influence of righ-censoring is limited. FHS records social ties across multiple dimensions (e.g., spouse, family, friends, neighbors). We define two individuals as connected if any type of social tie exists between them and include only ties formed prior to exam 2 to avoid endogeneity in network formation.
@@
Additional results from both applications are provided in Section C.3 of the Supplementary Material, including placebo tests supporting the identification assumptions, estimates based on a propensity score model that accounts for treatment diffusion, and estimated effects on blood pressure and cholesterol levels in the FHS.
@@ 7 Conclusion
This paper tackles the challenge of causal inference in longitudinal data with an unknown interference structure. We introduce a novel estimand, the AMR, to capture both direct and spillover effects generated by different treatment histories. This estimand is well-defined for any user-specified proximity metric, requiring only minimal knowledge of the underlying interference structure. It can be summarized by an MSM and identified under a slightly generalized form of sequential exchangeability. We show how to estimate any AMR and MSM parameters using a WLS estimator and establish its large-sample properties, including consistency and asymptotic normality, under mild restrictions on the dependence across units. The estimator reduces to the classic approach in longitudinal analysis when $d=0$, suggesting that its estimates have a causal interpretation under interference. We also develop HAC variance estimators for spillover effect estimates that are ensured to be asymptotically conservative. Our approach bridges methods in longitudinal data analysis and causal inference under interference, providing researchers with a powerful tool to study a broad range of causal effects in spatial and network settings, especially when interactions among units are poorly understood. We illustrate the method’s utility through simulations and two empirical applications, and discuss possible extensions.
Compared to alternative approaches based on outcome models or exposure mappings, our method avoids risks of model misspecification and the burden of computing exposure probabilities in settings with complex and unknown interference. However, it relies on researchers’ substantive understanding of the prevailing treatment assignment policy so that the propensity score is either known or can be consistently estimated. Future research should investigate more flexible strategies for propensity score estimation, such as sieve estimators or highly adaptive LASSO (Ertefaie, Hejazi and van der Laan, 2020). Another promising direction is to combine our approach with outcome modeling to improve robustness and efficiency. Finally, extending the framework to incorporate methods that account for right-censoring and binary outcomes would further enhance its applicability across biomedical and social science research.
@@
Appendix A Additional Technical Details
This section provides additional interpretations of the estimand and identification assumptions discussed in the main text, as well as technical details for the extensions introduced in Section 5.
@@ A.1 Structural Equations under the Identification Assumptions
Here, we consider $\mathcal{G}_{N}$ as fixed. $\varepsilon_{\mathbf{L}i}^{(t)}$, $\varepsilon_{Ai}^{(t)}$, and $\varepsilon_{Yi}$ represent other factors that may influence $\mathbf{L}_{i}^{(t)}$, $A_{i}^{(t)}$, and $Y_{i}$, respectively. They may include a unit’s own attributes, characteristics of its neighbors, and measurement errors. We do not assume that these error terms are identically distributed, nor do we impose independence on either $\varepsilon_{Yi}$ or $\varepsilon_{\mathbf{L}i}^{(t)}$.
@@
$\mathcal{N}_{\mathcal{G}_{N}}(i)$ in the subscripts represents a unit-specific and unknown neighborhood for unit $i$ in $\mathcal{G}_{N}$. The equations are compatible with the DAG in Figure 1 of the main text and closely resemble those in Ogburn et al. (2020). However, our framework allows for arbitrary heterogeneity in how units respond to treatment histories and does not impose a predefined exposure mapping that summarizes the effects of $\mathbf{A}_{\mathcal{N}}^{(1:T)}$. These equations formally define what unknown interference means and reflect a more realistic setting for spatial or network data, where both $Y_{i}$ and $\mathbf{L}_{i}^{(t)}$ may be shaped by unobserved and idiosyncratic factors, such as unit-specific neighborhood structures. This flexibility introduces the additional challenges for inference, as discussed in Section 4.
Let’s denote the collection $\left(\varepsilon_{\mathbf{L},\mathcal{i}}^{(1:T)},\varepsilon_{Y,\mathcal{i}}\right)$ as $\mathbf{F}_{i}^{(1:T)}$ and expectation over $\mathbf{F}_{i}^{(1:T)}$ as $\mathbb{E}[\cdot]$. Analyses in the main text are conditional on $\mathbf{F}_{i}^{(1:T)}$, with the expectation ${\rm E}\,[\cdot]={\rm E}\,\left[\cdot\mid\mathbf{F}_{i}^{(1:T)}\right]$. It is worth noting that ${\rm E}\,\left[Y_{i;j}\left(\mathbf{a}^{(1:T)}\right)\mid\mathbf{F}_{i}^{(1:T)},\mathbf{F}_{j}^{(1:T)}\right]={\rm E}\,\left[Y_{i;j}\left(\mathbf{a}^{(1:T)}\right)\mid\mathbf{F}_{i}^{(1:T)}\right]$ by definition. We can further marginalize $Y_{i;j}\left(\mathbf{a}^{(1:T)}\right)$ each AMR over $\mathbf{F}_{i}^{(1:T)}$ and obtain
$\mu^{*}\left(\mathbf{a}^{(1:T)};d\right)\coloneqq\frac{1}{N}\sum_{j=1}^{N}\mu_{j}\left(\Big{\{}\mathbb{E}\left[Y_{i;j}\left(\mathbf{a}^{(1:T)}\right)\right]\Big{\}}_{i\in\mathcal{N}};d\right),$ (11)
$\mu^{*}\left(\mathbf{a}^{(1:T)};d\right)=\mathbf{m}\left(\mathbf{a}^{(1:T)}\right)^{\prime}\boldsymbol{\beta}^{*}(d).\vskip-10.00002pt$ (12)
Below, we show that the main results hold for this estimand under additional restrictions.
A.2 Welfare Implication of the AMR
Consider the utilitarian welfare function: $W=\frac{1}{N}\sum_{i=1}^{N}{\rm E}\,\left[Y_{i}\left(\mathbf{A}_{\mathcal{N}}^{(1:T)}\right)\right]$, and assume that the treatment assignment policy in period $t$ is governed by a parameter $\gamma_{t}$: $P\left(A_{j}^{(t)}=1\mid\mathbf{V}_{j}^{(t)}\right)=g\left(\mathbf{V}_{j}^{(t)};\gamma_{t}\right)$. Then, we have
${\rm E}\,\left[Y_{i}\left(\mathbf{A}_{\mathcal{N}}^{(1:T)}\right)\right]={\rm E}\,\left[{\rm E}\,\left[Y_{i}\left(\mathbf{A}_{\mathcal{N}}^{(1:T)}\right)\mid\mathbf{V}_{j}^{(t)}\right]\right]$
$=$ ${\rm E}\,\left[\sum_{a_{j}^{(t)}\in\{0,1\}}{\rm E}\,\left[Y_{i}\left(a_{j}^{(t)},\mathbf{A}_{j}^{(1:(t-1),(t+1):T)},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right)\mid\mathbf{V}_{j}^{(t)},A_{j}^{(t)}=a_{j}^{(t)}\right]P\left(A_{j}^{(t)}=a_{j}^{(t)}\mid\mathbf{V}_{j}^{(t)}\right)\right].$
Using the fact that $P\left(A_{j}^{(t)}=0\mid\mathbf{V}_{j}^{(t)}\right)+P\left(A_{j}^{(t)}=1\mid\mathbf{V}_{j}^{(t)}\right)=1$, we know that
$\frac{\partial{\rm E}\,\left[Y_{i}\left(\mathbf{A}_{\mathcal{N}}^{(1:T)}\right)\right]}{\partial P\left(A_{j}^{(t)}=1\mid\mathbf{V}_{j}^{(t)}\right)}={\rm E}\,\left[Y_{i}\left(1,\mathbf{A}_{j}^{(1:(t-1),(t+1):T)},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right)\right]-{\rm E}\,\left[Y_{i}\left(0,\mathbf{A}_{j}^{(1:(t-1),(t+1):T)},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right)\right].$
Then, for any two period $t^{\prime}<t$, we can similarly obtain
$\frac{\partial^{T}{\rm E}\,\left[Y_{i}\left(\mathbf{A}_{\mathcal{N}}^{(1:T)}\right)\right]}{\partial P\left(\mathbf{A}_{j}^{(t^{\prime})}=1\mid\mathbf{V}_{j}^{(t^{\prime})}\right)\dots\partial P\left(\mathbf{A}_{j}^{(t)}=1\mid\mathbf{V}_{j}^{(t)}\right)}$
$=$ $\sum_{\mathbf{a}^{(t^{\prime}:t)}\in\mathcal{A}^{(s:t)}}(-1)^{t-\sum_{s=t^{\prime}}^{t}a^{(s)}}{\rm E}\,\left[Y_{i}\left(\mathbf{a}^{(t^{\prime}:t)},\mathbf{A}_{j}^{(1:t^{\prime},t:T)},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right)\right].$
Averaging units in the sample, we can see that
$\sum_{j=1}^{N}\frac{\partial^{T}W}{\partial P\left(\mathbf{A}_{j}^{(t^{\prime})}=1\mid\mathbf{V}_{j}^{(t^{\prime})}\right)\dots\partial P\left(\mathbf{A}_{j}^{(t)}=1\mid\mathbf{V}_{j}^{(t)}\right)}$
$=$ $\frac{1}{N}\sum_{j=1}^{N}\sum_{i=1}^{N}\sum_{\mathbf{a}^{(t^{\prime}:t)}\in\mathcal{A}^{(t^{\prime}:t)}}(-1)^{t-\sum_{s=t^{\prime}}^{t}a^{(s)}}{\rm E}\,\left[Y_{i}\left(\mathbf{a}^{(t^{\prime}:t)},\mathbf{A}_{j}^{(1:t^{\prime},t:T)},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right)\right]$
$=$ $\sum_{\mathbf{a}^{(t^{\prime}:t)}\in\mathcal{A}^{(t^{\prime}:t)}}(-1)^{t-\sum_{s=t^{\prime}}^{t}a^{(s)}}\sum_{d\in\mathcal{D}}\frac{1}{N}\sum_{j=1}^{N}\sum_{i=1}^{N}\mathbf{1}\{d_{ij}\in\Omega_{j}(d)\}{\rm E}\,\left[Y_{i}\left(\mathbf{a}^{(t^{\prime}:t)},\mathbf{A}_{j}^{(1:t^{\prime},t:T)},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right)\right]$
$=$ $\sum_{\mathbf{a}^{(t^{\prime}:t)}\in\mathcal{A}^{(t^{\prime}:t)}}(-1)^{t-\sum_{s=t^{\prime}}^{t}a^{(s)}}\sum_{d\in\mathcal{D}}\frac{1}{N}\sum_{j=1}^{N}\mu_{j}\left(\Big{\{}Y_{i;j}\left(\mathbf{a}^{(t^{\prime}:t)}\right)\Big{\}}_{i\in\mathcal{N}};d\right)|\Omega_{j}(d)|$
$\to$ $\sum_{\mathbf{a}^{(t^{\prime}:t)}\in\mathcal{A}^{(t^{\prime}:t)}}(-1)^{t-\sum_{s=t^{\prime}}^{t}a^{(s)}}\sum_{d\in\mathcal{D}}\mu\left(\mathbf{a}^{(t^{\prime}:t)};d\right)|\Omega(d)|.$
Finally, consider the policy parameters $(\gamma_{1},\dots,\gamma_{T})$, we have
$\sum_{j=1}^{N}\frac{\partial^{T}W}{\partial\gamma_{t^{\prime}}\dots\partial\gamma_{t}}=\sum_{j=1}^{N}\frac{\partial^{T}W}{\partial P\left(\mathbf{A}_{j}^{(t^{\prime})}=1\mid\mathbf{V}_{j}^{(t)}\right)\dots\partial P\left(\mathbf{A}_{j}^{(t)}=1\mid\mathbf{V}_{j}^{(t)}\right)}\prod_{s=t^{\prime}}^{t}\frac{\partial P\left(\mathbf{A}_{j}^{(s)}=1\mid\mathbf{V}_{j}^{(s)}\right)}{\partial\gamma_{t}}$
$=$ $\sum_{\mathbf{a}^{(t^{\prime}:t)}\in\mathcal{A}^{(t^{\prime}:t)}}(-1)^{t-\sum_{s=t^{\prime}}^{t}a^{(s)}}\sum_{d\in\mathcal{D}}\tilde{\mu}\left(\mathbf{a}^{(t^{\prime}:t)};d\right),$
where $\tilde{\mu}\left(\mathbf{a}^{(t^{\prime}:t)};d\right)=\frac{1}{N}\sum_{j=1}^{N}\mu_{j}\left(\Big{\{}Y_{i;j}\left(\mathbf{a}^{(t^{\prime}:t)}\right)\Big{\}}_{i\in\mathcal{N}};d\right)|\Omega_{j}(d)|\prod_{s=t^{\prime}}^{t}\frac{\partial P\left(\mathbf{A}_{j}^{(s)}=1\mid\mathbf{V}_{j}^{(s)}\right)}{\partial\gamma_{s}}$ is a re-weighted AMR conditional on the treatment history from period $t^{\prime}$ to period $t$. In particular, when $g\left(\mathbf{V}_{j}^{(t)};\gamma_{t}\right)=\gamma_{t}$ and $|\Omega_{j}(d)|=|\Omega(d)|+o(1)$, the expression above can be approximated by $\sum_{\mathbf{a}^{(t^{\prime}:t)}\in\mathcal{A}^{(t^{\prime}:t)}}(-1)^{t-\sum_{s=t^{\prime}}^{t}a^{(s)}}\sum_{d\in\mathcal{D}}\mu\left(\mathbf{a}^{(t^{\prime}:t)};d\right)|\Omega(d)|$. This result suggests that any partial derivative of the welfare function with regards to the policy parameters can be approximated by a linear combination of the (reweighted) sums of AMRs across proximity levels. Therefore, even though the AMRs do not identify optimal policies, they are informative about the directions of policy changes that improve welfare (Hu, Li and Wager, 2022; Viviano, 2020).
A.3 Technical Details of the Extensions
Placebo Tests
Figure A.1: A DAG illustration for placebo tests
$Y_{i}$$Y_{j}$$\mathbf{L}_{i}^{(2)}$$\mathbf{L}_{j}^{(2)}$$\mathbf{L}_{i}^{(1)}$$\mathbf{L}_{j}^{(1)}$$A_{i}^{(2)}$$A_{i}^{(1)}$$A_{j}^{(2)}$$A_{j}^{(1)}$$U_{i}^{(2)}$$U_{i}^{(1)}$$U_{i}^{(0)}$$U_{j}^{(2)}$$U_{j}^{(1)}$$U_{j}^{(0)}$
Diffusion and Dependence in Treatment Assignment
To allow for diffusion in treatment, we can modify Assumption 1 as follows:
Assumption A.1 (Sequential exchangeability with treatment diffusion)
$\Big{\{}Y_{i}\left(\mathbf{a}^{(1:T)},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right),\mathbf{L}_{i}^{(t+1)}\left(\mathbf{a}^{(1:T)},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right)\Big{\}}\perp A_{j}^{(t)}\mid\left(\mathbf{A}_{\{j\}\cup\mathcal{N}_{\mathcal{G}_{N}}(j,p)}^{(1:(t-1))},\mathbf{L}_{\{j\}\cup\mathcal{N}_{\mathcal{G}_{N}}(j,p)}^{(1:(t-1))}\right).\vskip-10.00002pt$
for any $i$, $j$, $t$, $\mathbf{a}^{(1:T)}$, and known $p$.
Here, $\mathcal{N}_{\mathcal{G}_{N}}(j,p)$ denotes the set of units whose proximity to $j$ does not exceed $p$, as defined in Section 4.1. Under Assumption A.1, the variables $\mathbf{A}_{\mathcal{N}_{\mathcal{G}_{N}}(j,p)}^{(1:(t-1))}$ and $\mathbf{L}_{\mathcal{N}_{\mathcal{G}_{N}}(j,p)}^{(1:(t-1))}$ must be included in $\mathbf{V}_{j}^{(t)}$ and controlled for when estimating unit $j$’s propensity score at time $t$. This adjustment is straightforward when $p$ is known and fixed, and it does not affect our theoretical results. We leave the case where $p$ is unknown or grows with $N$ to future work.
Assumption A.2 (Design mixing and design separation)
$\alpha_{INT,t}=o(N),\alpha_{EXT,t}=o(N),$
for any period $t$.
This assumption is satisfied by many common treatment assignment mechanisms, such as complete randomization. Formal proofs can be found in Sävje, Aronow and Hudgens (2021), where the authors also show that Assumption A.2 holds for common designs such as complete randomization.
Marginal Structural Models with a Large $T$
When $T$ is large, Assumption 4 may be questionable if effects from remote neighbors accumulate over time and eventually become non-negligible. To avoid this possibility, we denote an independent copy of the treatment history beyond the recent $s$ periods as $\mathbf{A}^{(1:T),s}_{\mathcal{N}_{N}}=\left(\mathbf{A}^{((T-s+1):T)}_{\mathcal{N}_{N}},\tilde{\mathbf{A}}^{(1:(T-s))}_{\mathcal{N}_{N}}\right)$. Then, Assumption 4 can be replaced by the following version:
Assumption A.3 (Approximate Neighborhood Interference in Two Dimensions)
$\max_{i\in\mathcal{N}_{N}}E\Bigg{|}Y_{i}\left(\mathbf{A}^{(1:T)}_{\mathcal{N}_{N}}\right)-Y_{i}\left(\mathbf{A}^{(1:T),s}_{\mathcal{N}_{N}}\right)\Bigg{|}\leq\theta_{T,s},$
$\max_{i\in\mathcal{N}_{N}}E\Bigg{|}Y_{i}\left(\mathbf{A}^{((T-s+1):T)}_{\mathcal{N}_{N}}\right)-Y_{i}\left(\mathbf{A}^{((T-s+1):T)}_{i,p}\right)\Bigg{|}\leq\theta_{N,p},$
where $\sup_{T}\theta_{T,s}\rightarrow 0$ as $s\rightarrow\infty$ and $\sup_{N}\theta_{N,p}\rightarrow 0$ as $p\rightarrow\infty$.
This assumption explicitly stipulates that the influence of remote observations, whether in the temporal or cross-unit dimension, will eventually be negligible in large samples. It is straightforward to verify that our main theorems hold if Assumption 4 is replaced by Assumption A.3.
Appendix B Proofs
This section provides the proofs of the theorems stated in the main text.
B.1 Proof for Proposition 1